OpenStudy (anonymous):

Prove that the inequality is true for all real numbers a,b,c:

OpenStudy (anonymous):

\[ 4(a^2 + b^2 + c^2) - \big[ (a+b)^2 + (b+c)^2 + (a+c)^2\big] \ge 0\]

OpenStudy (anonymous):

I got \( a^2 + b^2 + c^2 -ab - bc - ac \ge 0 \). But I don not know how to continue.

OpenStudy (anonymous):

I thought that I can somehow start with the fact that \((a+b+c)^2 \ge 0\).

OpenStudy (anonymous):

\[\Large a^2+b^2+c^2 \geq ab+bc+ac \] or \[\Large2a^2+2b^2+2c^2 \geq 2ab + 2bc + 2ac \]

OpenStudy (anonymous):

i would find the rate at which (a^2 + b^2 + c^2) incraeses and find the rate that (a+b)^2+(b+c)^2+(a+c)^2 increases if you know the first term increases faster, it will always be positive or >= 0

OpenStudy (anonymous):

\[\Large (a-b)^2+(b-c)^2+(c-a)^2 \geq 0  \]

OpenStudy (anonymous):

\[\Large a^2-2ab + b^2 + b^2 -2bc + c^2 + c^2 -2ac + a^2 \geq 0 \]

OpenStudy (anonymous):

So as in the expression above.

OpenStudy (anonymous):

The algebraic manipulation is a nightmare though, took me several tries too, I think I have seen something similar to this once a while ago, hence I remembered to add up the terms.

OpenStudy (anonymous):

Thanks Spacelimbus.

OpenStudy (anonymous):

You're welcome

OpenStudy (anonymous):

How did you get the idea of do that?

OpenStudy (anonymous):

See your original equation. Not the entire one, there it says. \[\Large (a+b)^2+(b+c)^2+(a+c)^2 = 2a^2 +2b^2 +2c^2 + 2ab + 2bc + 2ac \]

OpenStudy (anonymous):

After a while I thought that this is the same as we have in the step where you can say: \[\Large a^2+b^2+c^2 \geq ab + bc + ca \]

OpenStudy (anonymous):

So it's basically the same expression (after multiplication with 2) just that after algebraic manipulation the ab, bc, ca, terms will be negative, hence the minus sign in the binomials.